OpenStudy (anonymous):

a barbeque sells each stick for 25 pesos. if marginal equation is MC=300/X and the fixed cost is 5000 pesos. find the profit from selling 300 sticks.

OpenStudy (anonymous):

Profit = Revenue - Cost Profit = revenue per item x # of items sold - Cost Profit = (revenue per item)(items sold) - (Fixed Cost + Variable Cost) Profit = (25 pesos)(300 sticks) - (5000 pesos + 300/x) where x = items sold = 300, in this case.

OpenStudy (anonymous):

using indefinite integral?

OpenStudy (anonymous):

Profit = (unit price)(sales quantity) - (Fixed Cost - MC) MC is an equation involving x, which is also the sales quantity, which is 300 in this case. Sub in 300 for sales quantity and for the "x" in the MC equation, and solve... P = (25)(300) - (5000 - (300/300)) P = 7500 - (5000 - 1) P = 7500 - 4999 P = 2501 At least that's what I think...